OpenStudy (anonymous):

evaluate this integral ∫x^2 +1 /x^2+3x+2 dx

OpenStudy (lgbasallote):

have you tried dividing the polynomial?

OpenStudy (anonymous):

I did long division and got 1 -3x-1/x^2+3x+2

OpenStudy (lgbasallote):

you mean \[\int (1 - \frac{3x - 1}{x^2+3x+2})dx\] right?

OpenStudy (anonymous):

yea thats right

OpenStudy (lgbasallote):

split it up \[\int dx - \int \frac{3x-1}{x^2 + 3x +2}dx\] first one is simple second one is solveable by partial fractions

OpenStudy (anonymous):

oh yeah, i got it now thanks. I was actually trying to solve the second one using the substitution thats why i was getting the right answer.

OpenStudy (lgbasallote):

haha congrats :D